# Jet-Driven Formation of Bipolar Rings in Planetary Nebulae: Numerical Simulations Inspired by NGC 1514

### v1

@@ Abstract
We conduct three-dimensional hydrodynamical simulations of jets launched into a dense shell, reproducing two rings in a bipolar structure that resemble the two dusty rings of the planetary nebula (PN) NGC 1514. The scenario we simulate assumes that a strong binary interaction enhanced the mass loss rate from the asymptotic giant branch (AGB) stellar progenitor of NGC 1514, and shortly thereafter, the main-sequence companion accreted mass from the AGB star, launching a pair of jets. We find that adiabatic flows, where radiative losses are negligible, produce prominent rings, as observed in the infrared in NGC 1514. In contrast, when radiative cooling is significant, the rings are thin and faint. Our results reinforce the prevailing notion that jets play a substantial role in shaping planetary nebulae (PNe). More generally, as the binary companion to the central star of NGC 1514 avoided common envelope evolution, our results suggest that jets play a major role in many binary systems experiencing stable mass transfer at high rates. This conclusion complements the view that jets play a significant role in unstable mass transfer, specifically in common envelope evolution. Studies of strongly interacting binary systems, whether stable or not, should include jets. If jets continue to be active after ring formation, the outcomes are circum-jet rings, as observed in some other PNe and core-collapse supernova remnants.
@@ 2.1 The morphology of NGC 1514
Ressler et al. (2025) focused on the distinctive pair of rings, which is only visible in the IR and dominated by emission from small grain dust. In Section 3 we show that jets can form the rings’ morphology; these jets are earlier than those which shaped the inner structure of NGC 1514. Here, we first address the question of whether we can see signatures of departure from axisymmetry in the outer two rings that Ressler et al. (2025) analyzed. In panels b and d of Figure 1, we present two images from Ressler et al. (2025), a JWST IR image in $12.8\mu$m in panel b, and the rings with two ellipses (solid lines) that Ressler et al. (2025) drew, and one ellipse that we drew (dashed-red line). In panel b, we identify an arc. This shows that the outer dusty structure contains more than two smooth and symmetrical rings.
@@
Figure 2: A JWST $12.8\mu$m image adapted from Ressler et al. (2025) with our suggestion of a possibility of two bipolar rings that are much wider on the far side than on the near side.
@@ 2.2 A plausible scenario
In the second phase, the secondary star accretes mass and launches two opposite jets. The interaction is not purely axially symmetric due to several possible processes. ($i$) The dense shell that the AGB star ejected in the first phase might lack axial symmetry. ($ii$) The jet-launching companion is on an eccentric orbit. ($iii$) The jet-launching period is shorter, or not much longer, than the orbital period, as we suggest in Section 3.2.4. ($iv$) A tertiary object with an inclined orbit might influence the launching of the jets, particularly making them unequal. In Section 3, we present simulations of the interaction of the jets with the dense shell that the binary ejected in the first phase of the strong interaction. Our simulations in this initial study disregard the non-axially symmetric processes and the orbit of the jet-launching companion around the center of mass, as we aim only to demonstrate that jets can, in principle, form bipolar rings.
@@ 3.1 Numerical setup
We conduct three-dimensional (3D) hydrodynamical simulations using version 4.8 of the flash code (Fryxell et al., 2000), employing the unsplit hydro solver. FLASH is a modular, adaptive-mesh refinement (AMR) hydrodynamics and magnetohydrodynamics code.
[…]
@@
We set the initial radial velocity of the shell to be $v_{s}=10{~\rm km}{~\rm s}^{-1}$; this is a much slower velocity than the jets’ velocity, and its exact value is not of great significance to the goals of this study. We simulated cases with two different shells. One shell occupies the radial region $10^{17}{~\rm cm}<r<1.1\times 10^{17}{~\rm cm}$, with the density profile $\rho_{s}=1.6\times 10^{-19}(r/10^{17}{~\rm cm})^{-2}{~\rm g}{~\rm cm}^{-3}$, resulting in a total shell mass of $M_{s}=0.1M_{\odot}$. A mass loss episode with a mass loss rate of $\dot{M}_{s}=3.2\times 10^{-4}M_{\odot}{~\rm yr}^{-1}$ that lasted for $\Delta t_{s}=315{~\rm yr}$ formed this shell. The other shell we simulate occupies the radial region $10^{17}{~\rm cm}<r<1.2\times 10^{17}{~\rm cm}$, with the density profile $\rho_{s}=8\times 10^{-20}(r/10^{17}{~\rm cm})^{-2}{~\rm g}{~\rm cm}^{-3}$, resulting $M_{s}=0.1M_{\odot}$, $\dot{M}_{s}=1.6\times 10^{-4}M_{\odot}{~\rm yr}^{-1}$, and $\Delta t_{s}=630{~\rm yr}$.
@@
We inject two oppositely directed jets from a central region within a radius of $10^{16}{~\rm cm}$, aligned along the $z$-axis and confined within a half-opening angle of $\alpha_{j}=45^{\circ}$. We terminate the jet launching at $t=48{~\rm yrs}$. The jets are injected into the grid with a velocity of $v_{j}=800{~\rm km}{~\rm s}^{-1}$, slightly above the escape velocity from a main-sequence star. The combined mass-loss rate into both jets is $\dot{M}_{\rm 2jets}=1.1\times 10^{-4}M_{\odot}{~\rm yr}^{-1}$.
@@
In our simulations, we adopt an ideal-gas equation of state with a constant adiabatic index $\gamma$. To represent different thermodynamic conditions, we use two values: $\gamma=5/3$ and $\gamma=1.05$. The value $\gamma=5/3$ corresponds to a monoatomic ideal gas and is appropriate when radiative losses are negligible, representing nearly adiabatic evolution. To account for the effects of radiative cooling in an approximate manner, we also perform simulations with a significantly lower adiabatic index, $\gamma=1.05$. This low value mimics the cooling effect due to photon diffusion, allowing the gas to lose energy more efficiently and thereby approximating nearly isothermal behavior. This method has been commonly employed in astrophysical hydrodynamic simulations where explicit radiative transfer is computationally prohibitive (see our discussion in Akashi & Soker 2013). In addition, we also perform simulations with $\gamma=5/3$ but with radiative cooling of an optically thin plasma.
@@ 3.2.1 Density evolution: ring formation
Figure 3 shows the density maps in the meridional plane and at three times of the simulations with the initial wide dense shell and $\gamma=5/3$ with no radiative cooling; namely, an adiabatic simulation. At early times ($t=60~\mathrm{yr}$), the jets are still inside the dense shell and compressing material along the polar directions. After the jets break out from the dense shell, they inflate low-density polar bubbles and leave behind a dense barrel-like structure. The cross-section of the barrel-shaped 3D structure on the meridional plane is two arcs, as the middle panel of Figure 3 shows at $t=140~\mathrm{yr}$. The edges of these arcs are denser (contain a small red zone) than the rest of the arcs, and in 3D form a structure of two rings. This structure of two rings becomes more prominent at a later time of $t=220~\mathrm{yr}$, which we present in the right panel of Figure 3.
@@
Figure 4: Temperature maps in the same plane, the same times, and the same simulation as in Figure 3. Temperatures are expressed in Kelvin, as indicated by the color of the bar; note the varying upper values in the different panels ($4.6\times 10^{7}{~\rm K}$, $8.0\times 10^{7}{~\rm K}$, and $1.4\times 10^{7}{~\rm K}$. Note the hot bubble developing in the center.
In Figures 5 and 6 we present the density and temperature maps in the meridional plane of the simulation with an adiabatic index of $\gamma=1.05$, at the same times as in Figures 3 and 4. Other than the value of $\gamma$, the parameters of the two simulations are the same. The significantly lower value of $\gamma$ mimics a case where radiative cooling is substantial; therefore, the thermal energy in the shocked gas is substantially lower than in the adiabatic case. The interaction is momentum-driven rather than energy-driven (for studies on momentum- and energy-conserving flows, see, e.g., García-Segura & López 2000; Mellema 2001, 2003).
@@
Both simulations produce a barrel-shaped structure from the remnant of the dense shell, featuring two dense rings at the edges of the remnant. Both simulations inflate two bipolar lobes, but of different structures. In the two simulations, the jet-shell interaction in the polar directions is prone to Rayleigh-Taylor instabilities. This is evident from the clumpy polar shell in the middle panel of Figure 3 and two right panels of Figure 5.
The main differences between the adiabatic and the $\gamma=1.05$ simulations, as shown in the density and temperature maps, are as follows. (1) While the adiabatic interaction inflates two large polar bubbles expanding in all directions, as seen in the last temperature maps, the $\gamma=1.05$ accelerates material in the polar directions, as expected from a momentum-conserving flow. (2) The adiabatic simulation disperses the polar shell gas at early times, as no polar gas is seen in the right panel of Figure 3. Dense clumps are seen in the right panel of Figure 5. (3) At late times, the adiabatic simulation develops a hot central bubble, which the $\gamma=1.05$ does not. (4) The hot central bubble in the adiabatic simulation compresses the remnant of the dense shell. This does not occur in the $\gamma=1.05$ simulation.
Although both simulations produce a pair of bipolar rings, in Section 3.2.3 we demonstrate that the adiabatic one more closely matches the observed double-ring morphology of PN NGC 1514.
@@ 3.2.2 The flow structure
To better reveal the flow structure, we present in Figures 7 and 8 the velocity maps for the adiabatic and $\gamma=1.05$ simulations at the same time and plane as in Figures 3 and 5, respectively; the arrow present the direction of the flow and the color the velocity magnitude, from $10{~\rm km}{~\rm s}^{-1}$ (deep blue) to $500{~\rm km}{~\rm s}^{-1}$ (deep red). At early times, the fastest flows (green/yellow/red) are bipolar, corresponding to the launching directions of the jets. At later times, the two simulations differ.
@@
As expected, most of the fast-moving material in the momentum-conserving simulation of $\gamma=1.05$ (Figure 8) more or less maintains its bipolar outflow. In the adiabatic simulation, on the other hand, fast vortices develop around the center in the hot central bubble (Figure 7). These vortices play a role in compressing the barrel-shaped remnant of the dense shell (Figure 3). Additionally, the flow of the bipolar lobes in the adiabatic simulation, which is outside the rings (i.e., at larger distances from the center), exhibits lateral expansion, resulting in significant velocities at intermediate latitudes.
@@ 3.2.3 Mimicking observational images
In Figures 9 we present the column density, in units of ${~\rm g}{~\rm cm}^{-2}$ according to the color bar, at three times as in Figure 3, and for $i=50^{\circ}$. The coordinates are those on the plane of the sky, where $X_{S}$ is parallel to the numerical $X$-axis and $Z_{S}$ is in the YZ plane of the numerical grid. In Figure 10 we present the column density for the simulation with $\gamma=1.05$, at the same inclination angle and times as in Figure 9. Both simulations produce a bipolar double-ring system. However, while the adiabatic simulation $\gamma=5/3$ produces prominent thick rings, the $\gamma=1.05$ simulation produces thin rings that do not resemble the rings of NGC 1514.
@@
In Figure 11 we present the column density of the adiabatic case $\gamma=5/3$, where the initial AGB dense shell is thin, namely, half the initial thickness of the results we presented in previous figures. We present the column density map at the last time, $t=220{~\rm yr}$, but at three inclination angles, $i=25^{\circ}$ (upper panel; close to being pole-on), $i=50^{\circ}$ (middle panel), and $i=75^{\circ}$ (lower panel; close to being edge-on). The thin shell simulation also produces the bipolar double structure as in NGC 1514 for an inclination angle of $i\simeq 50^{\circ}$; we mark one ring with a black ellipse in the middle panel. The large inclination angle, almost edge-on, presents two rings, on the upper and lower parts of the simulated nebula. These two rings resemble the two rings of PN Abell 14 (e.g., Akras et al. 2016b for an image of Abell 14), although in Abell 14, there is a structure outside the two rings.
@@
Figure 12 compares the column density of the adiabatic simulation with a thin shell at $i=50^{\circ}$ that we present in the middle panel of Figure 11 with an IR observation of NGC 1514 (lower panel). We mark the two rings that we identify in this simulation by dashed-yellow ellipses (marked by a black ellipse in the middle panel of Figure 11). We place the two ellipses on the image of NGC 1514 in the lower panel. The simulation produces a structure of two bipolar rings that resembles that of NGC 1514. The simulated ring does not fit exactly the southeast ring. This is because of the departure from mirror symmetry that we discussed in Section 2 Overall, we conclude from this section that adiabatic jet-shell interactions can produce a bipolar pair of rings as observed in NGC 1514, and some other PNE.
Figure 12: A comparison of the column density of the adiabatic simulation with the thin shell at $i=50^{\circ}$ from the middle panel of Figure 11 to an IR image of NGC 1514 (panel d of Figure 1 adapted from Ressler et al. 2025, and rotated). We mark the rings produced by the simulated jets with yellow dotted ellipses and copied them to the IR image of NGC 1514 in the lower panel. The upper simulated ring fits the observed one, while there is a small departure in the fitting of the lower ring, resulting from departure from a mirror symmetry of NGC 1514 (Section 2).

### v2

@@ Abstract
We conduct three-dimensional hydrodynamical simulations of jets launched into a dense shell, reproducing two rings in a bipolar structure that resemble the two dusty rings of the planetary nebula (PN) NGC 1514. The scenario we simulate assumes that a strong binary interaction enhanced the mass loss rate from the asymptotic giant branch (AGB) stellar progenitor of NGC 1514, and shortly thereafter, the main-sequence companion accreted mass from the AGB star, launching a pair of jets. We find that flows with negligible radiative losses produce prominent rings, as observed in the infrared in NGC 1514. In contrast, when radiative cooling is significant, the rings are thin and faint. Our results reinforce the prevailing notion that jets play a substantial role in shaping planetary nebulae (PNe). More generally, as the binary companion to the central star of NGC 1514 avoided common envelope evolution, our results suggest that jets play a major role in many binary systems experiencing stable mass transfer at high rates. This conclusion complements the view that jets play a significant role in unstable mass transfer, specifically in common envelope evolution. Studies of strongly interacting binary systems, whether stable or not, should include jets. If jets continue to be active after ring formation, the outcomes are circum-jet rings, as observed in some other PNe and core-collapse supernova remnants.
@@ 2.1 The morphology of NGC 1514
[…]
@@
[…]
@@ 2.2 A plausible scenario
In the second phase, the secondary star accretes mass and launches two opposite jets. The interaction is not purely axially symmetric due to several possible processes. ($i$) The dense shell that the AGB star ejected in the first phase might lack axial symmetry. ($ii$) The jet-launching companion is on an eccentric orbit. (iiiiii) The jet-launching period is shorter, or not much longer, than the orbital period, as we suggest in Section 3.2.4. ($iv$) A tertiary object with an inclined orbit might influence the launching of the jets, particularly making them unequal. In Section 3, we present simulations of the interaction of the jets with the dense shell that the binary ejected in the first phase of the strong interaction. Our simulations in this initial study disregard the non-axially symmetric processes and the orbit of the jet-launching companion around the center of mass, as we aim only to demonstrate that jets can, in principle, form bipolar rings.
@@ 3.1 Numerical setup
We conduct three-dimensional (3D) hydrodynamical simulations using version 4.8 of the flash code (Fryxell et al., 2000), employing the unsplit hydro solver (Lee & Deane, 2009). FLASH is a modular, adaptive-mesh refinement (AMR) hydrodynamics and magnetohydrodynamics code. We do not include gravity, as the velocities of the gas are much larger than the escape velocity from the regions we simulate. In all simulations we used a Courant–Friedrichs–Lewy number of 0.3. For the Riemann solver we adopted the HLLC method (Toro et al., 1994), together with artificial viscosity to stabilize shocks. The computational domain is a Cartesian grid $(x,y,z)$ with outflow boundary conditions at all boundaries. The plane $z=0$ is the stellar and binary equatorial plane. Our simulations cover both sides of the equatorial plane. The total sizes of the grid in each direction are $L_{x}=Ly=4\times 10^{17}{~\rm cm}$, and $L_{z}=6\times 10^{17}{~\rm cm}$. We utilize a 3D AMR grid with eight refinement levels, resulting in the best resolution with a cell size of $2^{-10}L$, where $L$ is the grid size. The numerical code splits cells to reach higher resolution in zones where there are large gradients in one or more flow quantities. For the parameters we use, the grid cells are boxes, with the best resolution of $(\Delta_{x},\Delta_{y},\Delta_{z})=2^{-10}(L_{x},L_{y},L_{z})=(3.9\times 10^{14}{~\rm cm},3.9\times 10^{14}{~\rm cm},5.86\times 10^{14}{~\rm cm})$. We simulated two cases with a lower level of resolution, where the best resolution has twice as large cells. We still obtain the rings, but they are smeared somewhat, as expected. We therefore find that our simulations converge in the highest resolution we use. With our computer resources, we could not simulate a higher level of resolution.
@@
We set the initial radial velocity of the shell to be $v_{s}=10{~\rm km}{~\rm s}^{-1}$; this is a much slower velocity than the jets’ velocity, and its exact value is not of great significance to the goals of this study. We simulated cases with two different shells. One shell occupies the radial region $10^{17}{~\rm cm}<r<1.1\times 10^{17}{~\rm cm}$, with the density profile $\rho_{s}=1.6\times 10^{-19}(r/10^{17}{~\rm cm})^{-2}{~\rm g}{~\rm cm}^{-3}$, resulting in a total shell mass of $M_{s}=0.1M_{\odot}$. A mass loss episode with a mass loss rate of $\dot{M}_{s}=3.2\times 10^{-4}M_{\odot}{~\rm yr}^{-1}$ that lasted for $\Delta t_{s}=315{~\rm yr}$ formed this shell. The other shell we simulate occupies the radial region $10^{17}{~\rm cm}<r<1.2\times 10^{17}{~\rm cm}$, with the density profile $\rho_{s}=8\times 10^{-20}(r/10^{17}{~\rm cm})^{-2}{~\rm g}{~\rm cm}^{-3}$, resulting $M_{s}=0.1M_{\odot}$, $\dot{M}_{s}=1.6\times 10^{-4}M_{\odot}{~\rm yr}^{-1}$, and $\Delta t_{s}=630{~\rm yr}$. We tested the evolution of the shell when we do not inject jets. The shell expands because of the initial radial velocity, and widens because of internal pressure. What is important to the present study is that for the 220 years of the simulations we present here, the shell maintains its smooth and spherical structure. The instabilities we will report on are real and do not result from an unstable shell expansion.
@@
We numerically initialize the jets by defining two oppositely directed conical regions with a prescribed half-opening angle and injection radius. Within these cones, at each timestep during the jet active phase, the gas cells are assigned the jet velocity, density, and temperature that correspond to the desired jet properties (mass-loss rate, kinetic, and thermal power). The injection is maintained for the specified jet active time. This prescription ensures a continuous, steady injection of bipolar jets with well-defined parameters into the computational domain. In the present simulations, we inject two oppositely directed jets from a central region within a radius of $10^{16}{~\rm cm}$, aligned along the $z$-axis and confined within a half-opening angle of $\alpha_{j}=45^{\circ}$. We terminate the jet launching at $t=48{~\rm yrs}$. The jets are injected into the grid with a velocity of $v_{j}=800{~\rm km}{~\rm s}^{-1}$, slightly above the escape velocity from a main-sequence star. The combined mass-loss rate into both jets is $\dot{M}_{\rm 2jets}=1.1\times 10^{-4}M_{\odot}{~\rm yr}^{-1}$.
@@
In our simulations, we adopt an ideal gas equation of state with a uniform composition. Ionization is not included, and the gas is treated as neutral.
In our simulations, we adopt an ideal-gas equation of state with a constant adiabatic index $\gamma$, and a uniform composition of a neutral gas. We do not follow the ionization of the gas. We follow the global structure for which ionization will not have a large effect. As well, following ionization requires including the ionizing radiation of the central star, which is beyond the present scope. To represent different thermodynamic conditions, we use two values: $\gamma=5/3$ and $\gamma=1.05$. The value $\gamma=5/3$ corresponds to a monoatomic ideal gas and is appropriate when radiative losses are negligible (often referred to as "adiabatic evolution", even though entropy changes at shocks). To account for the effects of radiative cooling in an approximate manner, we also perform simulations with a significantly lower adiabatic index, $\gamma=1.05$. This low value mimics the cooling effect due to photon diffusion, allowing the gas to lose energy more efficiently and thereby approximating nearly isothermal behavior. This method has been commonly employed in astrophysical hydrodynamic simulations where explicit radiative transfer is computationally prohibitive (see our discussion in Akashi & Soker 2013).
In addition to $\gamma<5/3$ simulations, we also perform simulations with $\gamma=5/3$ but with radiative cooling of an optically thin plasma. In these simulations, we implement radiative cooling in FLASH by adding a tabulated cooling function from Sutherland & Dopita (1993a). The code interpolates the table during runtime to obtain the cooling function as a function of temperature. FLASH treats this cooling term in an operator-split manner: after the hydrodynamical update, the internal energy is updated separately according to the cooling rate. This procedure, standard in FLASH, ensures numerical stability even for short cooling timescales by using sub-cycling if needed. The method is first-order accurate in time with respect to the cooling source term, as is customary in operator-split implementations.
@@ 3.2.1 Density evolution: ring formation
Figure 3 shows the density maps in the meridional plane and at three times of the simulations with the initial wide dense shell and $\gamma=5/3$ with no radiative cooling. At early times ($t=60~\mathrm{yr}$), the jets are still inside the dense shell and compressing material along the polar directions. After the jets break out from the dense shell, they inflate low-density polar bubbles and leave behind a dense barrel-like structure. The cross-section of the barrel-shaped 3D structure on the meridional plane is two arcs, as the middle panel of Figure 3 shows at $t=140~\mathrm{yr}$. The edges of these arcs are denser (contain a small red zone) than the rest of the arcs, and in 3D form a structure of two rings. This structure of two rings becomes more prominent at a later time of $t=220~\mathrm{yr}$, which we present in the right panel of Figure 3.
@@
Figure 4: Temperature maps in the same plane, the same times, and the same simulation as in Figure 3. Temperatures are expressed in Kelvin, as indicated by the color of the bar; note the varying upper values in the different panels ($4.6\times 10^{7}{~\rm K}$, $8.0\times 10^{7}{~\rm K}$, and $1.4\times 10^{7}{~\rm K}$. Note the hot bubble developing in the center, and the filamentary temperature structure at late times that result from Rayleigh-Taylor instabilities.
In Figures 5 and 6 we present the density and temperature maps in the meridional plane of the simulation with an adiabatic index of $\gamma=1.05$, at the same times as in Figures 3 and 4. Apart from the value of $\gamma$, the parameters of the two simulations are identical. The significantly lower value of $\gamma$ mimics a case where radiative cooling is substantial; therefore, the thermal energy in the shocked gas is substantially lower than in the non-radiative case. The interaction is momentum-driven rather than energy-driven (for studies on momentum- and energy-conserving flows, see, e.g., García-Segura & López 2000; Mellema 2001, 2003).
@@
Both simulations produce a barrel-shaped structure from the remnant of the dense shell, featuring two dense rings at the edges of the remnant. Both simulations inflate two bipolar lobes, but of different structures. In the two simulations, the jet-shell interaction in the polar directions is prone to Rayleigh-Taylor instabilities. This is evident from the clumpy polar shell in the middle panel of Figure 3 and the two right panels of Figure 5. The temperature maps at late times (right panels in Figures 4 and 6), present filamentary structures inside the bubbles that result from Rayleigh-Taylor instabilities. We mapped (not shown) Rayleigh-Taylor unstable zones, and they also exhibit a filamentary structure, in addition to the instability at the front of the inflated bubble. These filaments are not along the grid axis and are curved. Therefore, these instabilities are Rayleigh-Taylor physical instabilities and not numerical instabilities.
The main differences between the non-radiative and the $\gamma=1.05$ simulations, as shown in the density and temperature maps, are as follows. (1) While the non-radiative interaction inflates two large polar bubbles expanding in all directions, as seen in the last temperature maps, the $\gamma=1.05$ accelerates material in the polar directions, as expected from a momentum-conserving flow. (2) The non-radiative simulation disperses the polar shell gas at early times, as no polar gas is seen in the right panel of Figure 3. Dense clumps are seen in the right panel of Figure 5. (3) At late times, the non-radiative simulation develops a hot central bubble, which the $\gamma=1.05$ does not. (4) The hot central bubble in thenon-radiative simulation compresses the remnant of the dense shell. This does not occur in the $\gamma=1.05$ simulation.
Although both simulations produce a pair of bipolar rings, in Section 3.2.3 we demonstrate that the non-radiative one more closely matches the observed double-ring morphology of PN NGC 1514.
@@ 3.2.2 The flow structure
To better reveal the flow structure, we present in Figures 7 and 8 the velocity maps for the non-radiative and $\gamma=1.05$ simulations at the same time and plane as in Figures 3 and 5, respectively; the arrow present the direction of the flow and the color the velocity magnitude, from $10{~\rm km}{~\rm s}^{-1}$ (deep blue) to $500{~\rm km}{~\rm s}^{-1}$ (deep red). At early times, the fastest flows (green/yellow/red) are bipolar, corresponding to the launching directions of the jets. At later times, the two simulations differ.
@@
As expected, most of the fast-moving material in the momentum-conserving simulation of $\gamma=1.05$ (Figure 8) more or less maintains its bipolar outflow. In the non-radiative simulation, on the other hand, fast vortices develop around the center in the hot central bubble (Figure 7). These vortices play a role in compressing the barrel-shaped remnant of the dense shell (Figure 3). Additionally, the flow of the bipolar lobes in the non-radiative simulation, which is outside the rings (i.e., at larger distances from the center), exhibits lateral expansion, resulting in significant velocities at intermediate latitudes.
@@ 3.2.3 Mimicking observational images
In Figures 9 we present the column density, in units of ${~\rm g}{~\rm cm}^{-2}$ according to the color bar, at three times as in Figure 3, and for $i=50^{\circ}$. The coordinates are those on the plane of the sky, where $X_{S}$ is parallel to the numerical $X$-axis and $Z_{S}$ is in the YZ plane of the numerical grid. In Figure 10 we present the column density for the simulation with $\gamma=1.05$, at the same inclination angle and times as in Figure 9. Both simulations produce a bipolar double-ring system. However, while the non-radiative simulation $\gamma=5/3$ produces prominent thick rings, the $\gamma=1.05$ simulation produces thin rings that do not resemble the rings of NGC 1514.
@@
In Figure 11 we present the column density of the non-radiative case $\gamma=5/3$, where the initial AGB dense shell is thin, namely, half the initial thickness of the results we presented in previous figures. We present the column density map at the last time, $t=220{~\rm yr}$, but at three inclination angles, $i=25^{\circ}$ (upper panel; close to being pole-on), $i=50^{\circ}$ (middle panel), and $i=75^{\circ}$ (lower panel; close to being edge-on). The thin shell simulation also produces the bipolar double structure as in NGC 1514 for an inclination angle of $i\simeq 50^{\circ}$; we mark one ring with a black ellipse in the middle panel. The large inclination angle, almost edge-on, presents two rings, on the upper and lower parts of the simulated nebula. These two rings resemble the two rings of PN Abell 14 (e.g., Akras et al. 2016b for an image of Abell 14), although in Abell 14, there is a structure outside the two rings.
@@
Figure 12 compares the column density of the non-radiative simulation with a thin shell at $i=50^{\circ}$ that we present in the middle panel of Figure 11 with an IR observation of NGC 1514 (lower panel). We mark the two rings that we identify in this simulation by dashed-yellow ellipses (marked by a black ellipse in the middle panel of Figure 11). We place the two ellipses on the image of NGC 1514 in the lower panel. The simulation produces a structure of two bipolar rings that resembles that of NGC 1514. The simulated ring does not fit exactly the southeast ring. This is because of the departure from mirror symmetry that we discussed in Section 2 Overall, we conclude from this section that non-radiative jet-shell interactions can produce a bipolar pair of rings as observed in NGC 1514, and some other PNE.
Figure 12: A comparison of the column density of the non-radiative simulation with the thin shell at $i=50^{\circ}$ from the middle panel of Figure 11 to an IR image of NGC 1514 (panel d of Figure 1 adapted from Ressler et al. 2025, and rotated). We mark the rings produced by the simulated jets with yellow dotted ellipses and copied them to the IR image of NGC 1514 in the lower panel. The upper simulated ring fits the observed one, while there is a small departure in the fitting of the lower ring, resulting from departure from a mirror symmetry of NGC 1514 (Section 2).